OpenStudy (vegan555):

Draw 150° in standard position. Then find a if the point (a, 1) is on the terminal side of 150°.

OpenStudy (shamim):

150°=90°+60° Right?

OpenStudy (shamim):

150=90+60 Right?

OpenStudy (shamim):

So its in second quadrunt?

OpenStudy (vegan555):

yes

OpenStudy (vegan555):

would it be -root 3/2, and 1?

OpenStudy (shamim):

180-30=150 degree?

OpenStudy (shamim):

Anyway drawing is urgent for better understanding

OpenStudy (shamim):

But i m frm android phone nd unable to draw it

OpenStudy (shamim):

I want to say that Tan 30=-1/a Right?

OpenStudy (vegan555):



OpenStudy (shamim):

Yes

OpenStudy (vegan555):